### v2

@@ 3.4 Reward design
While our primary learning signal is derived from unsupervised game outcomes to minimize reward engineering, we incorporate two auxiliary rewards to ensure stable training. The final reward signal is composed of three components:
@@
  • Action format regularization: To ensure parsable outputs, we provide a small positive reward (+0.05) for validly formatted actions and a large negative penalty (-10.0) that terminates the game for invalid ones, similar to the approach in DeepSeek-R1 (guo2025deepseek).
  • Response length penalty: To encourage conciseness, we apply a turn-level penalty for verbosity, inspired by Kimi k1.5 (team2025kimi), which scales linearly for responses longer than a set threshold. The penalty is calculated as:
    $r_{\text{length}}(l)=\alpha\cdot\max\left(0,1-\frac{l-l_{\text{min}}}{l_{\text{max}}-l_{\text{min}}}\right)$ (4)
@@ 4 Experiments
In this section, we present a series of extensive experiments designed to validate MARSHAL’s ability to obtain generalizable multi-agent skills. Additional results can be found in the Appendix.
@@ 4.1 Experimental setup
We use Qwen-4B (yang2025qwen3), a state-of-the-art instruction-tuned model, as our foundation to measure the incremental multi-agent capabilities unlocked by MARSHAL. We train two model types: specialist agents on single games (Tic-Tac-Toe, Kuhn Poker, Mini Hanabi) and a generalist agent on all three simultaneously. Our primary baseline is SPIRAL, a recent work focused on self-play in purely competitive, zero-sum games (liu2025spiral). Our evaluation is structured in four stages:
@@
    Strategic ability: We first evaluate strategic competence by benchmarking agents against strong Monte Carlo Tree Search (MCTS) or Nash Equilibrium (NE) opponents. Performance is measured over 1000 games on both the training games and a suite of more complex, held-out variants. For the cooperative Hanabi games, we report the standard self-play score.
@@
    Generalization to multi-agent systems: This is the primary test of this work. We integrate MARSHAL agents into two established systems: the competitive MAD framework (liang2023encouraging) and the cooperative AutoGen framework (wu2024autogen), measuring performance on downstream math and QA benchmarks.
@@ 4.2 Strategic ability
Our analysis begins with strategic ability on games, with normalized game return detailed in Fig 3. The MARSHAL framework proves highly effective: specialist agents not only master their training domains and outperform baselines, but also generalize effectively to more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). We also observe evidence of cross-category skill transfer. For example, the Tic-Tac-Toe specialist model demonstrates smooth improvement not only on Tic-Tac-Toe, but also the OOD Mini Hanabi (Fig. 4), suggesting MARSHAL cultivates foundational skills like turn-based planning that are broadly beneficial.
@@ 4.3 Generalization to multi-agent systems
The ultimate test of MARSHAL is whether skills honed in games transfer to practical, out-of-domain challenges. We evaluate this on a suite of demanding mathematics and general QA benchmarks, including MATH500 (cobbe2021training), GSM8K, AQUA-RAT (ling2017program), AIME24, AMC23, MMLU-STEM (hendrycks2020measuring), and GPQA-Diamond (rein2024gpqa). As a preliminary step, we investigate how our game tasks enhances foundational reasoning in a standard single-agent setting. Notably, MARSHAL models achieve notable improvements over both the Qwen-4B baseline on a number of math benchmarks, on par with SPIRAL (Table 1).
We further embed our agents into established multi-agent systems to directly measure their cooperative and competitive capabilities. In the competitive MAD debate framework, agents trained on competitive games show a clear advantage. Notably, the generalist agent is able to achieve an average of 3.51% over the original Qwen3-4B across all benchmarks. Conversely, in the cooperative AutoGen framework, the agents trained for cooperation—the Hanabi specialist and the generalist—excel across numerous benchmarks. In particular, the generalist model achieves a striking gain of 6.57% on GPQA-Diamond in the MAD framework, and 10.00% on AIME in the AutoGen framework.
These results provide compelling evidence that MARSHAL forges distinct, transferable skills for both competition and collaboration. These capabilities directly translate to improved performance in downstream multi-agent applications, with the generalist model proving to be the most robust.
@@ 4.4 Reasoning pattern and failure mode analysis
To understand the mechanisms behind MARSHAL’s successful generalization, we analyze the agent’s reasoning (<think>) traces. This qualitative analysis reveals the emergence of key multi-agent skills cultivated in our mix of strategic games. We highlight two representative patterns corresponding to the competitive and cooperative training environments, as showcased in Table 2.
First, reflecting the skills honed in competitive games, MARSHAL develops a role-aware strategy. In Tic-Tac-Toe, the agent explicitly identifies itself as the "second-move player" and adopts a defensive strategy. This skill transfers directly to the MAD debate framework, where the agent recognizes its role as the "negative side assistant" and adopts a negative stance rather than just solving the problem.
More profoundly, MARSHAL cultivates intent recognition, a key component of Theory of Mind. In the cooperative Hanabi, the agent learns to interpret a teammate’s actions as communications with hidden intent (e.g., "Maybe they want me to play this card?"). This sophisticated skill transfers to the AutoGen framework, where the agent infers a collaborator’s uncertainty from subtle cues, such as a missing TERMINATE token, instead of treating it as a simple error. Together, these patterns provide qualitative evidence that MARSHAL goes beyond improving benchmark scores; it equips the agent with the cognitive toolkit essential for effective multi-agent interaction.
To further substantiate these qualitative observations with quantitative evidence, we performed a failure mode analysis on the GPQA-Diamond benchmark within the MAD framework. Adopting the taxonomy of cemri2025multi, we categorized failures into System Design Issues (e.g., formatting errors, loop repetition), Inter-Agent Misalignment (e.g., multi-agent reasoning failures like ignoring peers or task derailment), and Task Verification issues.
As shown in Fig 5 (left), while MARSHAL improves basic instruction following (reducing System Design Issues by $\sim$7%), the reduction in Inter-Agent Misalignment is significantly larger (11.5%). To understand the drivers of this strategic improvement, we further decompose the Inter-Agent Misalignment category into its sub-categories. The breakdown reveals that the performance gains are primarily driven by reductions in Task Derailment and Ignored Other Agent’s Input. This indicates that MARSHAL agent are actively listening to peers and maintaining focus on its objective, validating our hypothesis that game-theoretic self-play cultivates transferable multi-agent reasoning skills.
@@ 4.5 Ablation studies
Training agents against fixed, expert opponents reveals a intuitive and critical flaw: the agents overfit. As shown in Table 3, these models fail to generalize, exhibiting significant performance drop on games outside their direct training domain. This is particularly evident with the Kuhn Poker specialist, which suffers a severe degradation on non-poker games—a clear case of strategic mode collapse. These results confirm that the adaptive curriculum provided by self-play is essential for the development of robust, generalizable policies.
@@
Tic-Tac-Toe Kuhn Poker
Mini
Hanabi
Connect Four
Leduc
Hold’em
Simple
Hanabi
@@
Tic-Tac-Toe Kuhn Poker
Mini
Hanabi
Connect Four
Leduc
Hold’em
Simple
Hanabi
@@ 5 Related work
To further extend the ability of LLMs with collaborative intelligence, LLM-based multi-agent systems have been proposed for various tasks. AutoGen (wu2024autogen) and CAMEL (li2023camel) designs cooperative agents for general reasoning and question answering. MetaGPT (hong2024metagpt) and ChatDev (qian2023chatdev) propose static multi-agent workflows with specialized roles for software development tasks. Multi-Agent Debate (liang2023encouraging) takes a competitive approach by letting LLMs propose and criticize solutions over multiple rounds for the final answer. These works mainly focus on designing workflows for multi-agent interactions with fixed LLMs. Other work have proposed to generate workflows automatically (zhang2024aflow). Our work takes a complementary approach by training LLMs to enhance their multi-agent capabilities and build stronger foundation models that can be integrated with these systems.
Reinforcement learning for LLMs.
@@
Reinforcement learning (RL) has emerged as a prominent approach to enhance LLMs’ ability from instruction following to reasoning. To align LLMs with human values, RL from human feedback (RLHF) (ouyang2022training) and from AI feedback (RLAIF) (bai2022constitutional) have been widely used in post-training to steer models towards favorable behaviors. The recent success of reasoning models (jaech2024openai; guo2025deepseek) further popularize RL from verifiable rewards (RLVR) for incentivizing long Chain-of-Thought reasoning capability of LLMs and have been applied in mathematical reasoning (shao2024deepseekmath), coding (wei2025swe), and tool-using (jin2025search). However, these works primarily consider single-turn or single-agent tasks, while our approach extends RL training to multi-turn, multi-agent scenarios. Recently, MT-GRPO (zeng2025reinforcing) also address the turn-level credit assignment issue for multi-turn RL. However, they propose to normalize the turn-level rewards turn-by-turn, which can introduce variance in strategic games with varying number of turns. We propose a "sum-then-normalize" approach that takes advantage of all turns in the normalization, increasing robustness.
Self-play training of LLMs.
@@
The great success of achieving superhuman performance in multi-agent games (silver2016mastering; vinyals2019grandmaster; berner2019dota) has spurred the use of self-play to train LLMs by playing with themselves (zhang2024survey). Some work (chen2024self; wu2024self) combines game-theoretic property of self-play to overcome intransitivity of human preference in LLM alignment. Another line of work (meta2022human; xu2023language; xu2025learning) uses self-play to improve the strategic ability of LLM agents in specific games. The most relevant work to ours is SPAG (cheng2024self) and SPIRAL (liu2025spiral), which also train LLMs via self-play to show generalization to reasoning tasks. However, these work focuses on zero-sum game and show generalization in single-agent evaluations. In contrast, our work consider both cooperative and competitive games and demonstrate generalized improvement in multi-agent system. As a concurrent work to ours, SPIRAL also address the heterogeneity of player role in advantage estimation and introduces a Role-Conditioned Advantage Estimation (RAE), analogous to our agent-specific normalization. We advance this finding by revealing that role separation is only critical in games with distinct return distributions between players.
@@ 6 Discussion
In this work, we introduced MARSHAL, a framework utilizing self-play in a diverse range cooperative and competitive strategic games to cultivate multi-agent capabilities in LLMs. Our findings demonstrate that skills honed in strategic games directly transfer to enhanced performance in general multi-agent systems, establishing self-play as a scalable paradigm for training LLM agents.
Despite these promising results, limitations remain. First, our study utilizes two-player games as efficient "prototypes" to cultivate foundational skills. While we demonstrate that these constrained settings are sufficient for strong transfer to general tasks, scaling to larger $N$-player environments introduces significantly greater challenges regarding non-stationarity, population diversity, and credit assignment that warrant dedicated future investigation. Additionally, moving beyond classic games to complex "social sandboxes" (e.g., simulated software engineering, collaborative researching) represents a compelling direction for training agents in more practical domains.
@@
This work adheres to the ICLR Code of Ethics. The primary ethical consideration of our work is the dual-use nature of enhancing LLM agent capabilities. While our goal is to foster beneficial multi-agent skills like cooperation and competition, the resulting models could potentially be applied to malicious ends. We release our code and models to the research community to encourage further study into the safety and alignment of more capable agents.
@@ Appendix B Implementation details
Our implementation of MARSHAL is built upon ROLL (wang2025reinforcement), a robust open-source codebase for post-training LLMs with reinforcement learning. ROLL’s native support for agentic, multi-turn rollouts provided a strong foundation for our system. To achieve high performance, ROLL leverages vLLM for efficient inference during the rollout phase (kwon2023efficient) and is built on Megatron-LM for distributed training (megatron-lm). All game environments used in this work were implemented using OpenSpiel (LanctotEtAl2019OpenSpiel) and VS-Bench (xu2025vs), ensuring the correctness and standardization of game logic.
@@
We use Qwen-4B as our base model for all experiments reported in our main results. The training is conducted in a fully online manner, where self-play trajectories are immediately used for policy updates. Specialist agents are trained with a batch size of 128 trajectories, while the generalist agent uses a batch size of 384 (128 from each game). All models were trained for a total of 200 optimization steps. For optimization, we employ the Adam optimizer (kingma2014adam) with a cosine annealing learning rate schedule, which warms up for 10 steps to a peak of $1\times 10^{-6}$ before gradually decaying to 0.
@@
Appendix C Evaluation details
Strategic ability.
  • Kuhn Poker: The exact Nash Equilibrium (NE) policy.
Generalization to multi-agent systems.
For downstream evaluations, we assess all models on a comprehensive suite of mathematics and general QA benchmarks. In the single-agent setting, we utilize evaluation scripts from Qwen2.5-Math (yang2024qwen2; yang2024qwen2_5_math) for MATH500, GSM8K, AQUA-RAT, AIME24, AMC23, and MMLU-STEM, while employing lm-evaluation-harness (eval-harness) for GPQA-Diamond. For multi-agent evaluations, we adopt the MASLab framework (ye2025maslab) across all benchmarks. To balance evaluation efficiency and accuracy, we set the maximum output length to 8,192 tokens for all settings.
Appendix D Scaling to larger models
Strategic ability.
Qwen3-8B 48.38 33.12 27.00 10.48 7.26 4.55
MARSHAL (Generalist, 8B) 54.05 44.49 55.28 21.55 53.89 37.27
Generalization to multi-agent systems.
MAS Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MAD Qwen3-8B 82.49 95.00 96.36 83.46 70.00 90.00 89.59 53.03
AutoGen Qwen3-8B 79.68 88.80 95.91 83.07 60.00 89.19 89.30 51.52
Appendix E Comparison with SPIRAL: decoupling algorithm and game environments
  1. 1.
  2. 2.
E.1 Strategic ability comparison
Ablation on algorithm.
Ablation on games.
When training only on competitive games (row 3), the agent naturally fails to acquire cooperative skills (low scores in Hanabi). This serves as a crucial control baseline for analyzing the downstream transfer results below.
MARSHAL (Full Method) 53.13 40.05 54.33 19.98 54.56 36.75
SPIRAL (RAE) + Our Games 48.35 42.05 24.93 9.70 6.11 5.53
MARSHAL Alg. + Adv. Games 48.80 47.55 24.48 17.90 29.98 11.48
E.2 Generalization to multi-agent systems
We further evaluate how these variations affect transfer to general reasoning benchmarks within multi-agent systems.
Generalization to competitive systems (MAD).
Generalization to cooperative systems (AutoGen).
Table 9 reveals the critical role of the game environments. The agent trained on the competitive-only games suffers a significant performance drop in the cooperative AutoGen framework (Average 80.34% vs. 82.15% for MARSHAL). This confirms that cooperative games are not merely "more data"; they are a necessary component for transferring skills to cooperative downstream tasks. Additionally, the RAE-based agent also underperforms, again solidifying the novelty of the algorithmic design of MARSHAL.
Table 8: Generalization to the competitive MAD framework.
Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MARSHAL (Full Method) 75.96 92.80 95.60 83.86 46.67 80.00 87.36 45.45
Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MARSHAL (Full Method) 82.15 95.20 94.54 86.61 66.67 92.50 89.53 50.00
Conclusion.
Appendix F Comparison with MT-GRPO
A concurrent work MT-GRPO (zeng2025reinforcing) also identifies the limitations of trajectory-level estimation in the original GRPO and proposes a fine-grained turn-level strategy. While MT-GRPO shares our motivation, and is also a critical first step towards fine-grain credit assignment in multi-turn GRPO, our "sum-then-normalize" approach differs fundamentally from their strategy, offering distinct advantages in stability and scalability in the game tasks that we consider in this work.
F.1 Theoretical differences
  1. 1.
  2. 2.
F.2 Empirical Comparison
Strategic ability.
Generalization to MAS.
We further extended this comparison to the integration within multi-agent systems to assess if the stability issues in game training affect general reasoning transfer. Table 11 present the results in the competitive MAD and cooperative AutoGen frameworks, respectively. MARSHAL consistently achieves higher average scores (75.01% vs. 73.77% in MAD; 80.15% vs. 79.10% in AutoGen) and outperforms MT-GRPO on key hard reasoning benchmarks like AIME and AMC. This indicates that the training proficiency provided by MARSHAL translates directly to better robust reasoning in general multi-agent tasks.
MARSHAL (Ours) 53.70 38.79 50.48 22.75 43.00 29.75
MT-GRPO 50.10 40.05 36.67 18.55 39.94 20.08
MAD MARSHAL (Ours) 75.01 92.20 96.06 83.07 43.33 82.50 86.76 41.12
MT-GRPO 73.77 92.60 95.91 84.65 36.67 77.50 86.91 42.13
MT-GRPO 79.10 95.40 94.62 85.04 56.67 90.00 89.02 42.93
Appendix G Hyperparameter analysis on response length penalty
Selection of length constraints.
Ablation on penalty weight $\alpha$.
Model Setup Strategic Performance Response Statistics
Tic-Tac-
Toe
Kuhn
Poker
Mini
Hanabi
Connect
Four
Leduc
Hold’em
Simple
Hanabi
Avg Len
OL in
Mini
Hanabi
OL in
Simple
Hanabi
Analysis.
The results yield two key conclusions:
  1. 1.
  2. 2.
Appendix H Hyperparameters

### v3

@@ 3.4 Reward design
While our primary learning signal is derived from the intrinsic game outcomes to minimize reward engineering, we incorporate two auxiliary rewards to ensure stable training. The final reward signal is composed of three components:
@@
  • Format reward: To ensure parsable outputs, we provide a small positive reward (+0.05) for validly formatted actions and a large negative penalty (-10.0) that terminates the game for invalid ones, similar to the approach in DeepSeek-R1 (Guo et al., 2025).
  • Response length penalty: To encourage conciseness, we apply a turn-level penalty for verbosity, inspired by Kimi k1.5 (Team et al., 2025), which scales linearly for responses longer than a set threshold. The penalty is calculated as
    $r_{\text{length}}(l)=\alpha\cdot\max\left(0,1-\frac{l-l_{\text{min}}}{l_{\text{max}}-l_{\text{min}}}\right),$ (4)
@@ 4 Experiments
In this section, we present extensive experiment results to validate MARSHAL’s ability to obtain generalizable multi-agent reasoning capability. Additional results can be found in the Appendix.
@@ 4.1 Experimental setup
We use Qwen-4B (Yang et al., 2025), a state-of-the-art reasoning model, as our foundation to measure the multi-agent reasoning capabilities of MARSHAL. We train two model types: specialist agents on each game (Tic-Tac-Toe, Kuhn Poker, Mini Hanabi) and a generalist agent on all three simultaneously. Implementation details and hyperparameter settings are provided in Appendix B and C. Our primary baseline is SPIRAL, a recent work focused on self-play in purely competitive, zero-sum games (Liu et al., 2025). Our evaluation is structured in four stages (see Appendix D for details):
@@
    Strategic ability: We first evaluate strategic competence by benchmarking agents against strong Monte Carlo Tree Search (MCTS) or Nash Equilibrium (NE) opponents. Performance is measured over 1000 games on both the training games and a suite of more complex, held-out games. For the cooperative Hanabi games, we report the standard self-play score.
@@
    Generalization to multi-agent systems: This is the primary test of this work. We integrate MARSHAL agents into two established systems: the competitive MAD framework (Liang et al., 2023) and the cooperative AutoGen framework (Wu et al., 2024a), measuring zero-shot generalization on standard math and QA benchmarks.
@@ 4.2 Strategic ability
Our analysis begins with strategic ability on games, with normalized game return detailed in Fig 3. The MARSHAL framework proves highly effective: specialist agents not only master their training domains and outperform baselines, but also generalize effectively to more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). We also observe evidence of cross-category skill generalization. For example, the Tic-Tac-Toe specialist model demonstrates smooth improvement not only on Tic-Tac-Toe, but also the OOD Mini Hanabi (Fig. 4), suggesting MARSHAL cultivates foundational skills like turn-based planning that are broadly beneficial.
@@ 4.3 Generalization to multi-agent systems
The ultimate test of MARSHAL is whether skills honed in games generalize to practical, out-of-domain challenges. We evaluate this on a suite of demanding mathematics and QA benchmarks in a zero-shot manner, including MATH500 (Cobbe et al., 2021), GSM8K, AQUA-RAT (Ling et al., 2017), AIME24, AMC23, MMLU-STEM (Hendrycks et al., 2020), and GPQA-Diamond (Rein et al., 2024). As a preliminary step, we investigate how our game tasks enhance foundational reasoning in a standard single-agent setting. Notably, MARSHAL models achieve notable improvements over both the Qwen-4B baseline on a number of math benchmarks, on par with SPIRAL (Table 1).
We further embed our agents into established multi-agent systems to directly measure their cooperative and competitive capabilities. In the competitive MAD debate framework, agents trained on competitive games show a clear advantage. Notably, the generalist agent is able to achieve an average of 3.51% over the original Qwen3-4B across all benchmarks. Conversely, in the cooperative AutoGen framework, the agents trained for cooperation—the Hanabi specialist and the generalist—excel across numerous benchmarks. In particular, the generalist model achieves a striking gain of 7.57% on GPQA-Diamond in the MAD framework, and 10.00% on AIME in the AutoGen framework.
These results provide compelling evidence that MARSHAL forges distinct, generalizable skills for both competition and collaboration. These capabilities directly translate to improved performance in downstream multi-agent applications, with the generalist model proving to be the most robust.
@@ 4.4 Reasoning pattern and failure mode analysis
To understand the mechanisms behind MARSHAL’s successful generalization, we analyze the agent’s reasoning (<think>) traces. This qualitative analysis reveals the emergence of key multi-agent reasoning patterns cultivated in our mix of strategic games. We highlight two representative patterns corresponding to the competitive and cooperative training environments, as showcased in Table 2.
First, reflecting the skills honed in competitive games, MARSHAL develops a role-aware strategy. In Tic-Tac-Toe, the agent explicitly identifies itself as the "second-move player" and adopts a defensive strategy. This skill generalizes directly to the MAD debate framework, where the agent recognizes its role as the "negative side assistant" and adopts a negative stance rather than just solving the problem.
More profoundly, MARSHAL cultivates intent recognition, a key component of Theory of Mind. In the cooperative Hanabi, the agent learns to interpret a teammate’s actions as communications with hidden intent (e.g., "Maybe they want me to play this card?"). This sophisticated skill generalizes to the AutoGen framework, where the agent infers a collaborator’s uncertainty from subtle cues, such as a missing TERMINATE token, instead of treating it as a simple error. Together, these patterns provide qualitative evidence that MARSHAL goes beyond improving benchmark scores; it equips the agent with the cognitive toolkit essential for effective multi-agent interaction.
To further substantiate these qualitative observations with quantitative evidence, we performed a failure mode analysis on the GPQA-Diamond benchmark within the MAD framework. Adopting the taxonomy of Cemri et al. (2025), we categorized failures into System Design Issues (e.g., formatting errors, loop repetition), Inter-Agent Misalignment (e.g., multi-agent reasoning failures like ignoring peers or task derailment), and Task Verification issues.
As shown in Fig 5 (left), while MARSHAL improves basic instruction following (reducing System Design Issues by $\sim$7%), the reduction in Inter-Agent Misalignment is significantly larger (11.5%). To understand the drivers of this strategic improvement, we further decompose the Inter-Agent Misalignment category into its sub-categories. The breakdown reveals that the performance gains are primarily driven by reductions in Task Derailment and Ignored Other Agent’s Input. This indicates that MARSHAL agent are actively listening to peers and maintaining focus on its objective, validating our hypothesis that game-theoretic self-play cultivates generalizable multi-agent reasoning skills.
@@ 4.5 Ablation studies
Training agents against fixed, expert opponents reveals an intuitive and critical flaw: the agents overfit. As shown in Table 3, these models fail to generalize, exhibitinga significant performance drop on games outside their direct training domain. This is particularly evident with the Kuhn Poker specialist, which suffers a severe degradation on non-poker games—a clear case of strategic mode collapse. These results confirm that the adaptive curriculum provided by self-play is essential for the development of robust, generalizable policies.
@@
Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
@@
Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
@@ 5 Related work
LLM-based multi-agent systems. To further extend the ability of LLMs with collaborative intelligence, LLM-based multi-agent systems have been proposed for various tasks. AutoGen (Wu et al., 2024a) and CAMEL (Li et al., 2023) design cooperative agents for general reasoning and question answering. MetaGPT (Hong et al., 2024) and ChatDev (Qian et al., 2023) propose static multi-agent workflows with specialized roles for software development tasks. Multi-Agent Debate (MAD) (Liang et al., 2023) takes a competitive approach by letting LLMs propose and criticize solutions over multiple rounds for the final answer. These works mainly focus on designing workflows for multi-agent interactions with fixed LLMs. Other works have proposed to generate workflows automatically (Zhang et al., 2024a). In comparison, our work takes a complementary approach by training LLMs to enhance their multi-agent reasoning capabilities and build stronger foundation models that can be integrated with these frameworks.
@@
Reinforcement learning for LLMs. Reinforcement learning (RL) has emerged as a prominent approach to enhance LLMs’ ability from instruction following to reasoning. To align LLMs with human values, RL from human feedback (RLHF) (Ouyang et al., 2022) and from AI feedback (RLAIF) (Bai et al., 2022) have been widely used in post-training to steer models towards favorable behaviors. The recent success of reasoning models (Jaech et al., 2024; Guo et al., 2025) further popularizes RL from verifiable rewards (RLVR) for incentivizing long Chain-of-Thought reasoning capability of LLMs and has been applied in mathematical reasoning (Shao et al., 2024), coding (Wei et al., 2025), and tool-using (Jin et al., 2025). However, these works primarily consider single-turn or single-agent tasks, while our approach extends RL training to multi-turn, multi-agent scenarios. Recently, MT-GRPO (Zeng et al., 2025) also address the turn-level credit assignment issue for multi-turn RL. However, they propose to normalize the turn-level rewards turn-by-turn, which can introduce variance in strategic games with varying numbers of turns. We propose a "sum-then-normalize" approach that takes advantage of all turns in the normalization, increasing robustness.
@@
Self-play training of LLMs. The great success of achieving superhuman performance in multi-agent games (Silver et al., 2016; Vinyals et al., 2019; Berner et al., 2019) has spurred the use of self-play to train LLMs by playing with themselves (Zhang et al., 2024b). Some work (Chen et al., 2024; Wu et al., 2024b) combines the game-theoretic property of self-play to overcome the intransitivity of human preference in LLM alignment. Another line of work (FAIR et al., 2022; Xu et al., 2023; 2025a) uses self-play to improve the strategic ability of LLM agents in specific games. The most relevant work to ours is SPAG (Cheng et al., 2024) and SPIRAL (Liu et al., 2025), which also train LLMs via self-play to show generalization to reasoning tasks. However, this work focuses on zero-sum games and considers generalization in single-agent evaluations. In contrast, our work consider both cooperative and competitive games and demonstrate generalized improvement in multi-agent systems. As a concurrent work to ours, SPIRAL also addresses the heterogeneity of player roles in advantage estimation and introduces a Role-Conditioned Advantage Estimation (RAE), analogous to our agent-specific normalization. We advance this finding by revealing that role separation is only critical in games with distinct return distributions between players.
@@ 6 Discussion
In this work, we introduced MARSHAL, a framework utilizing self-play in a diverse range of cooperative and competitive strategic games to cultivate multi-agent reasoning capabilities in LLMs. Our findings demonstrate that skills honed in strategic games directly translate to enhanced performance in general multi-agent systems, establishing self-play as a scalable paradigm for training LLM agents.
Despite these promising results, limitations remain. First, our study utilizes two-player games as efficient "prototypes" to cultivate foundational skills. While we demonstrate that these constrained settings are sufficient for strong generalization to standard reasoning tasks, scaling to larger $N$-player environments introduces significantly greater challenges regarding non-stationarity, population diversity, and credit assignment that warrant dedicated future investigation. Additionally, moving beyond classic games to complex "social sandboxes" (e.g., simulated software engineering, collaborative research) represents a compelling direction for training agents in more practical domains.
@@
This work adheres to the ICLR Code of Ethics. The primary ethical consideration of our work is the dual-use nature of enhancing LLM agent capabilities. While our goal is to foster beneficial multi-agent reasoning skills like cooperation and competition, the resulting models could potentially be applied to malicious ends. We release our code and models to the research community to encourage further study into the safety and alignment of more capable agents.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No.62406159, 62325405), Ant Group, Beijing National Research Center for Information Science and Technology (BNRist), Beijing Innovation Center for Future Chips, Shenzhen Key Laboratory of Ubiquitous Data Enabling (No. ZDSYS20220527171406015), and Tsinghua Shenzhen International Graduate School-Shenzhen Pengrui Endowed Professorship Scheme of Shenzhen Pengrui Foundation.
@@ Appendix B Implementation details
Our implementation of MARSHAL is built upon ROLL (Wang et al., 2025), a robust open-source codebase for post-training LLMs with reinforcement learning. ROLL’s native support for agentic, multi-turn rollouts provided a strong foundation for our system. To achieve high performance, ROLL leverages vLLM for efficient inference during the rollout phase (Kwon et al., 2023) and is built on Megatron-LM for distributed training (Shoeybi et al., 2019). All game environments used in this work were implemented using OpenSpiel (Lanctot et al., 2019) and VS-Bench (Xu et al., 2025b), ensuring the correctness and standardization of game logic.
@@
We use Qwen-4B as our base model for all experiments reported in our main results. The training is conducted in a fully online manner, where self-play trajectories are immediately used for policy updates. Specialist agents are trained with a batch size of 128 trajectories, while the generalist agent uses a batch size of 384 (128 from each game). All models were trained for a total of 200 optimization steps. For optimization, we employ the Adam optimizer (Kingma, 2014) with a cosine annealing learning rate schedule, which warms up for 10 steps to a peak of $1\times 10^{-6}$ before gradually decaying to 0.
@@
Appendix C Hyperparameters